# Realising VCD for untwisted automorphism groups of RAAGs

### v1

@@ 2.4. Culler–Vogtmann Outer space $CV_{n}$ and its spine $K_{n}$
For surveys of some of the applications of $CV_{n}$, we refer the reader to [Vog02], [Vog18], [Vog].
Generalising $CV_{n}$, Charney and Vogtmann (initially with Crisp [CCV07], then with Stambaugh [CSV17], and finally the full construction with Bregman [BCV23]) have produced an ‘Outer space’ $\mathcal{O}_{\Gamma}$ for any RAAG $A_{\Gamma}$. In particular, there is an untwisted Outer space $\Sigma_{\Gamma}$, first constructed in [CSV17], which is a contractible complex with a proper action of $U(A_{\Gamma})$. In analogy with $CV_{n}$ retracting onto its spine $K_{n}$, untwisted Outer space also has a spine, denoted $K_{\Gamma}$. We will describe this in more detail in §3.2.
@@ 2.5. Virtual cohomological dimension
Charney and Vogtmann have shown [CV09] that for any finite simplicial graph $\Gamma$$\operatorname{Out}(A_{\Gamma})$ is virtually torsion-free and has finite virtual cohomological dimension. Hence $\textsc{vcd}(U(A_{\Gamma}))$ is also finite.
We have the following geometric control over the virtual cohomological dimension.
@@
Since the spine $K_{n}$ of $CV_{n}$ is proper (since it is a locally finite complex) and contractible and has a proper and cocompact action of $\operatorname{Out}(F_{n})$, this theorem implies that $\textsc{vcd}(\operatorname{Out}(F_{n}))\leq\dim(K_{n})=2n-3$. On the other hand, as noted in §0 of [CV86], one can find a free abelian subgroup of $\operatorname{Out}(F_{n})$ of rank $32n-$ 3; one such example is the image in $\operatorname{Out}(F_{n})$ of
[…]
@@ 3. The untwisted spine and known results on $\textsc{vcd}(U(A_{\Gamma}))$
For any right-angled Artin group $A_{\Gamma}$, untwisted Outer space $\Sigma_{\Gamma}$ is built analogously. The Salvetti complex $\mathbb{S}_{\Gamma}$ generalises the $n$-petalled rose, and then the basic objects of $\Sigma_{\Gamma}$ are marked metric $\Gamma$-complexes: cube complexes which are ‘marked’ with a certain homotopy equivalence to $\mathbb{S}_{\Gamma}$ and which are endowed with a metric that turns the constituent cubes into rectilinear parallelepipeds. Once again, the untwisted spine $K_{\Gamma}$ is roughly obtained by ignoring this metric. We will not need the construction of the untwisted Outer space $\Sigma_{\Gamma}$, but we now outline in some detail the structure of $K_{\Gamma}$.
@@ 3.1. $\Gamma$-Whitehead partitions
We will now define $\Gamma$-Whitehead partitions. Roughly, given a vertex $m\in V$, we distribute the connected components of $\Gamma^{\pm}\setminus\operatorname{lk}({m})^{\pm}$ between two ‘sides’, with $m$ in one of the sides and $m^{-1}$ in the other.
@@
Let $m\in V$. A $\Gamma$-Whitehead partition based at m $(\mathcal{P},m)$ is a partition of the set $V^{\pm}$ that satisfies the following conditions:
@@
  • $\operatorname{lk}({\mathcal{P}})$ consists of all generators that commute with $m$, and their inverses—that is, $\operatorname{lk}({\mathcal{P}})=V\left(\operatorname{lk}({v})^{\pm}\right)$;
@@
We will abbreviate ‘$\Gamma$-Whitehead partition’ to ‘$\Gamma$-partition’; when $\Gamma$ is clear, we will often simply say ‘partition’ instead.
@@
[…]
@@
$\operatorname{split}({\mathcal{P}})\coloneqq\{v\in V\;:\;\mathcal{P}\;\text{splits}\;v\}\;;$
$\operatorname{max}({\mathcal{P}})\coloneqq\{v\in V\;:\;v\;\text{maximal w.r.t.~{}$\leq$ in}\;\operatorname{split}({\mathcal{P}})\}.$
The definition of $\Gamma$-partition is motivated by the following: each $\Gamma$-partition $(\mathcal{P},m)$ corresponds to a $\Gamma$-Whitehead automorphism $\varphi(\mathcal{P},m)$ of $A_{\Gamma}$, as follows. If $y\in\operatorname{split}({\mathcal{P}})$ then $\varphi(\mathcal{P},m)$ sends $y\mapsto ym^{-1}$ for $y\in P$, and $y\mapsto my$ if $y\in\overline{P}$. If $y,y^{-1}\in P$ then $\varphi(\mathcal{P},m)$ sends $y\mapsto mym^{-1}$, and for all other $y$, we have $\varphi(\mathcal{P},m)(y)=y$. Observe that if we did not require thickness (that is, that both sides contain at least two elements), then the corresponding $\Gamma$-automorphism might be inner.
@@
Say that $\Gamma$-partitions $\mathcal{P}$ and $\mathcal{Q}$ are adjacent if $\operatorname{max}({\mathcal{P}})\subseteq\operatorname{lk}({\mathcal{Q}})$, or equivalently, if $\operatorname{max}({\mathcal{Q}})\subseteq\operatorname{lk}({\mathcal{P}})$. (This equivalence follows from the pairwise equivalence of elements of $\operatorname{max}({\mathcal{P}})$.) If $v$ is a vertex of $\Gamma$, say that $v$ is adjacent to $\mathcal{P}$ if $v\in\operatorname{lk}({\mathcal{P}})$. Finally, say that two vertices of $\Gamma$ are adjacent if they are adjacent in $\Gamma$ in the usual graph-theoretic sense.
If two $\Gamma$-partitions $\mathcal{P},\mathcal{Q}$ are adjacent, then this means that some (hence any) base of $\mathcal{P}$ commutes with some (hence any) base of $\mathcal{Q}$, but that these two bases are not equal. We refer the reader to the “Warning” recorded after Definition 2.7 in [BCV23].
@@
Note that compatibility is not transitive; see Figure 5.
@@
[…]
@@ 3.2. The untwisted spine $K_{\Gamma}$
Let $H$ be a hyperplane in a cube complex $X$. Call the closure of the set of cubes that intersect $H$ the hyperplane carrier $\kappa(H)$. We say that $H$ is a carrier retract if $\kappa(H)$ is isomorphic to $H\times[0,1]$ as a cube complex. If $H$ is a carrier retract, then the hyperplane collapse associated with $H$ is the deformation retract $c_{H}\colon X\to X\smash{\sslash}H$ which collapses $\kappa(H)$ orthogonally to $H$. For a collection $\mathcal{H}$ of hyperplanes which are carrier retracts, we write $X\smash{\sslash}\mathcal{H}$ for the space obtained from $X$ by collapsing all the hyperplanes in $\mathcal{H}$ in any order. We will say that such a set $\mathcal{H}$ is acyclic if $X\to X\smash{\sslash}\mathcal{H}$ is a homotopy equivalence. [The double slash notation $X\smash{\sslash}\mathcal{H}$ is not intended to evoke other notions using the double slash, such as homotopy quotients or GIT quotients; we take our lead from e.g. [BCV24].]
@@
    If $\Gamma$ has $n$ vertices and no edges, then $\mathbb{S}_{\Gamma}^{\Pi}$ will be a finite connected graph with no separating edges or bivalent vertices and with fundamental group $A_{\Gamma}\cong F_{n}$. The edges dual to the set of hyperplanes $\{H_{\mathcal{P}}\}_{\mathcal{P}\in\Pi}$ form a maximal tree $T$, and the edges in $\mathbb{S}_{\Gamma}^{\Pi}\setminus T$ are oriented and labelled by $V(\Gamma)$. Collapsing the maximal tree $T$ yields a rose with $n$ petals, each labelled by a generator of $A_{\Gamma}\cong F_{n}$—this is a copy of the Salvetti complex.
@@
    The blowup $\mathbb{S}_{\Gamma}^{\mathcal{P}}$ of the Salvetti complex by a single partition $\mathcal{P}$ has the following structure. It has two vertices, say $x_{1}$ and $x_{2}$. The carrier of $H_{\mathcal{P}}$ is isomorphic to the product of an interval with the Salvetti complex for $\operatorname{lk}({\mathcal{P}})$. Let $v\in V(\Gamma)$. If there are two edges dual to $H_{v}$, then $v\in\operatorname{lk}({\mathcal{P}})$. If there is only one edge dual to the (oriented) hyperplane $H_{v}$ and it terminates at $P_{i}$, then $v\in P_{i}$; if the edge originates at $x_{i}$ then $v^{-1}\in P_{i}$. In this way one can recover $\mathcal{P}=\left(P_{1}\;|\;P_{2}\;|\;\operatorname{lk}({\mathcal{P}})\right)$ from the blowup $\mathbb{S}_{\Gamma}^{\mathcal{P}}$.
@@
If $\alpha$ is an untwisted marking, we call the equivalence class of the pair $(X,\alpha)$ a marked $\Gamma$-complex. If $X\cong\mathbb{S}_{\Gamma}$, we call the equivalence class of $(X,\alpha)$ a marked Salvetti. In this paper, ‘marked $\Gamma$-complex’ will always mean ‘untwisted marked $\Gamma$-complex’.
@@
Suppose that $\Pi=\{\mathcal{P}\}$. For each $m\in\operatorname{max}({\mathcal{P}})$, there is a cube complex isomorphism of $\mathbb{S}_{\Gamma}^{\mathcal{P}}$. Letting $c_{\mathcal{P}}^{-1}$ be a homotopy inverse for $c_{\mathcal{P}}$, the composition $\smash{c_{\mathcal{P}}\circ h_{m}\circ c_{\mathcal{P}}^{-1}\colon\mathbb{S}_{\Gamma}\to\mathbb{S}_{\Gamma}^{\mathcal{P}}\xrightarrow{\cong}\mathbb{S}_{\Gamma}^{\mathcal{P}}\to\mathbb{S}_{\Gamma}}$ induces the $\Gamma$-Whitehead automorphism $\varphi(\mathcal{P},m)$ on $A_{\Gamma}=~{}\pi_{1}\left(\mathbb{S}_{\Gamma}\right)$.
A set $\mathcal{T}$ of hyperplanes in a $\Gamma$-complex is called treelike if the collapse $c_{\mathcal{T}}\colon X\to X\smash{\sslash}\mathcal{T}$ gives a cube complex isomorphic to $\mathbb{S}_{\Gamma}$.
@@
Hence the star of a marked Salvetti $(\mathbb{S}_{\Gamma},\alpha)$ is a cube complex, with one $k$-cube for each compatible set $\Pi$ such that $|\Pi|=k$. We will denote this cube by specifying its two extreme vertices and the marking with which we view it, as $c\left(\emptyset,\Pi;\alpha\right)$. Faces of this cube correspond to subsets $\Pi_{1}$$\Pi_{2}$ with $\emptyset\subseteq\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi$, and will be denoted $c\left(\Pi_{1},\Pi_{2};\alpha\right)$. We suppress the marking from the notation when it is not needed. Note that a cube $c\left(\Pi_{1},\Pi_{2}\right)$ may lie in the star of two different marked Salvettis $(\mathbb{S}_{\Gamma},\alpha)$ and $(\mathbb{S}_{\Gamma},\beta)$. In this case, writing $c\left(\Pi_{1},\Pi_{2};\alpha\right)$ simply conveys that we are presently thinking of this cube as a subface of a cube in the star of $(\mathbb{S}_{\Gamma},\alpha)$.
@@
Charney–Stambaugh–Vogtmann [CSV17] also build a full untwisted Outer space $\Sigma_{\Gamma}$ for $U(A_{\Gamma})$, in analogy with Culler–Vogtmann Outer space $CV_{n}$ for the free group $F_{n}$. The spine $K_{\Gamma}$ naturally sits inside $\Sigma_{\Gamma}$ as a deformation retract, generalising the way $K_{n}$, the spine of $CV_{n}$, sits inside reduced Culler–Vogtmann Outer space $CV_{n}^{red}$ (the deformation retract of $CV_{n}$ obtained by restricting only to graphs with no separating edges). However, we will only be interested in the intrinsic structure of $K_{\Gamma}$ as the geometric realisation of a poset.
@@
As an immediate corollary, applying Theorem 2.7, we have

### v2

@@ 2.4. Culler–Vogtmann Outer space $CV_{n}$ and its spine $K_{n}$
For surveys of some of the applications of $CV_{n}$, we refer the reader to [VOG02], [VOG18], [CHV25].
Generalising $CV_{n}$, Charney and Vogtmann (initially with Crisp [CCV07], then with Stambaugh [CSV17], and finally the full construction with Bregman [BCV23]) have produced an ‘Outer space’ $\mathcal{O}_{\Gamma}$ for any RAAG $A_{\Gamma}$. In particular, there is an untwisted Outer space $\mathcal{O}_{\Gamma}^{U}$, first constructed in [CSV17] (where it is denoted $\Sigma_{\Gamma}$), which is a contractible complex with a proper action of $U(A_{\Gamma})$. In analogy with $CV_{n}$ retracting onto its spine $K_{n}$, untwisted Outer space also has a spine, denoted $K_{\Gamma}$. We will describe this in more detail in §3.2.
@@ 2.5. Virtual cohomological dimension
[…]
The following geometric control over the virtual cohomological dimension will be crucial.
@@
Since the spine $K_{n}$ of $CV_{n}$ is proper (since it is a locally finite complex) and contractible and has a proper and cocompact action of $\operatorname{Out}(F_{n})$, this theorem implies that
$\textsc{vcd}(\operatorname{Out}(F_{n}))\leq\dim(K_{n})=2n-3.$
On the other hand, as noted in §0 of [CV86], one can find a free abelian subgroup of $\operatorname{Out}(F_{n})$ of rank 2n32n-3; one such example is the image in $\operatorname{Out}(F_{n})$ of
$\langle\alpha_{i}:x_{i}\mapsto x_{1}x_{i},\;\beta_{i}:x_{i}\mapsto x_{i}x_{1}\;:\;2\leq i\leq n\rangle\leq\operatorname{Aut}(F_{n}).$
@@ 3. The untwisted spine and known results on $\textsc{vcd}(U(A_{\Gamma}))$
For any right-angled Artin group $A_{\Gamma}$, untwisted Outer space $\mathcal{O}_{\Gamma}^{U}$ is built analogously. The Salvetti complex $\mathbb{S}_{\Gamma}$ generalises the $n$-petalled rose, and then the basic objects of $\mathcal{O}_{\Gamma}^{U}$ are marked metric $\Gamma$-complexes: cube complexes which are ‘marked’ with a certain homotopy equivalence to $\mathbb{S}_{\Gamma}$ and which are endowed with a metric that turns the constituent cubes into rectilinear parallelepipeds. Once again, the untwisted spine $K_{\Gamma}$ is roughly obtained by ignoring this metric. We will not need the construction of the untwisted Outer space $\mathcal{O}_{\Gamma}^{U}$, but we now outline in some detail the structure of $K_{\Gamma}$.
@@ 3.1. $\Gamma$-Whitehead partitions
We now define $\Gamma$-Whitehead partitions. Roughly, given a vertex $m\in V$, we distribute the connected components of $\Gamma^{\pm}\setminus\operatorname{lk}({m})^{\pm}$ between two ‘sides’, with $m$ in one of the sides and $m^{-1}$ in the other.
@@
Let $m\in V$. A $\Gamma$-Whitehead partition $(\mathcal{P},m)$ based at m is a partition of the set $V^{\pm}$ satisfying the following conditions:
@@
  • $\operatorname{lk}({\mathcal{P}})=V\left(\operatorname{lk}({v})^{\pm}\right)$ – that is, $\operatorname{lk}({\mathcal{P}})$ consists of all generators that commute with $m$, and their inverses;
@@
We will abbreviate $\Gamma$-Whitehead partition’ to $\Gamma$-partition’; when $\Gamma$ is clear, we will often simply say ‘partition’ instead.
@@
Figure 4. This is a schematic showing an example of a graph $\Gamma$ and a $\Gamma$-partition $\mathcal{P}=\left(\{v,x,C_{2}^{\pm}\}\;|\;\{v^{-1},x^{-1},y,y^{-1},C_{1}^{\pm}\}\;|\;\operatorname{lk}({\mathcal{P}})\right)$, which is based at $v$. In the illustration of $\mathcal{P}$, the nodes are arranged in pairs, with a vertex of $\Gamma$ on the top row and its inverse below it on the bottom row. Here, $v$ is the only possible base for $\mathcal{P}$. The connected components of $\Gamma^{\pm}\setminus\operatorname{st}({v})^{\pm}$ are the singletons $\{x\}$$\{x^{-1}\}$$\{y\}$$\{y^{-1}\}$, along with the components $C_{1}^{\pm}$$C_{2}^{\pm}$ as indicated in the diagram.
@@
$\operatorname{split}({\mathcal{P}})\coloneqq\{v\in V\colon\mathcal{P}\;\text{splits}\;v\}\;$;
$\operatorname{max}({\mathcal{P}})\coloneqq\{v\in V\colon v\;\text{maximal w.r.t.~$\leq$ in}\;\operatorname{split}({\mathcal{P}})\}.$
The definition of $\Gamma$-partition is motivated by the following: each $\Gamma$-partition $(\mathcal{P},m)$ corresponds to a $\Gamma$-Whitehead automorphism $\varphi(\mathcal{P},m)$ of $A_{\Gamma}$, as follows. If $y\in\operatorname{split}({\mathcal{P}})$ then $\varphi(\mathcal{P},m)$ sends $y\mapsto ym^{-1}$ for $y\in P$, and $y\mapsto my$ if $y\in\overline{P}$. If $y,y^{-1}\in P$ then $\varphi(\mathcal{P},m)$ sends $y\mapsto mym^{-1}$, and for all other $y$, we have $\varphi(\mathcal{P},m)(y)=y$. Observe that if we did not require thickness (that is, that both sides contain at least two elements), then the corresponding $\Gamma$-Whitehead automorphism might be inner.
@@
Say that $\Gamma$-partitions $\mathcal{P}$ and $\mathcal{Q}$ are adjacent if $\operatorname{max}({\mathcal{P}})\subseteq\operatorname{lk}({\mathcal{Q}})$, or equivalently, if $\operatorname{max}({\mathcal{Q}})\subseteq\operatorname{lk}({\mathcal{P}})$ (equivalence follows from the pairwise equivalence of elements of $\operatorname{max}({\mathcal{P}})$.) If $v$ is a vertex of $\Gamma$, say that $v$ is adjacent to $\mathcal{P}$ if $v\in\operatorname{lk}({\mathcal{P}})$. Finally, say that two vertices of $\Gamma$ are adjacent if they are adjacent in $\Gamma$ in the usual graph-theoretic sense.
If two $\Gamma$-partitions $\mathcal{P},\mathcal{Q}$ are adjacent, then this means that some (hence any) base of $\mathcal{P}$ commutes with some (hence any) base of $\mathcal{Q}$, but that these two bases are not equal. The converse also holds. We refer the reader to the “Warning” recorded after Definition 2.7 in [BCV23].
@@
Note that compatibility is not transitive, as demonstrated in Figure 5.
@@
Figure 5. An example of a graph $\Gamma$ and three $\Gamma$-partitions, $\mathcal{P}_{1}$ (based at $a$), $\mathcal{P}_{2}$ (based at $b$), and $\mathcal{P}_{3}$ (based at $d$). Observe that although $e$ is split by $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$, it cannot serve as a base for either as $\operatorname{lk}({e})=\emptyset$, so $e<_{\circ}b$ and $e<_{\circ}d$$\mathcal{P}_{1}$ is compatible with $\mathcal{P}_{2}$ since there is a choice of side of each (shaded) which have empty intersection (in fact, $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are also adjacent, since $a$ and $b$ commute). Since $d$ does not commute with $a$ or $b$, we cannot have $d$ in the link of either $\mathcal{P}_{1}$ or $\mathcal{P}_{2}$, so $\mathcal{P}_{3}$ is not adjacent to $\mathcal{P}_{1}$ or $\mathcal{P}_{2}$. However, $P_{2}\cap\overline{P_{3}}=\emptyset$, so $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$ are compatible. Each side of $\mathcal{P}_{3}$ has non-empty intersection with each side of $\mathcal{P}_{1}$; hence $\mathcal{P}_{1}$ is not compatible with $\mathcal{P}_{3}$. This example shows that compatibility of partitions is not transitive.
@@ 3.2. The untwisted spine $K_{\Gamma}$
Let $H$ be a hyperplane in a cube complex $X$. Call the closure of the set of cubes that intersect $H$ the hyperplane carrier $\kappa(H)$. We say that $H$ is a carrier retract if $\kappa(H)$ is isomorphic to $H\times[0,1]$ as a cube complex. If $H$ is a carrier retract, then the hyperplane collapse associated with $H$ is the deformation retract $c_{H}\colon X\to X\smash{\sslash}H$ which collapses $\kappa(H)$ orthogonally to $H$. For a collection $\mathcal{H}$ of carrier retracts, we write $X\smash{\sslash}\mathcal{H}$ for the space obtained from $X$ by collapsing all the hyperplanes in $\mathcal{H}$ in any order (by [CSV17, Lemma 4.4], this is well-defined). We will say that such a set $\mathcal{H}$ is acyclic if $X\to X\smash{\sslash}\mathcal{H}$ is a homotopy equivalence. [The double slash notation $X\smash{\sslash}\mathcal{H}$ is not intended to evoke other notions using the double slash, such as homotopy quotients or GIT quotients; we take our lead from e.g. [BCV25].]
@@
    If $\Gamma$ has $n$ vertices and no edges, then $\mathbb{S}_{\Gamma}^{\Pi}$ will be a finite connected graph with no separating edges or bivalent vertices and with fundamental group $A_{\Gamma}\cong F_{n}$. The edges dual to the set of hyperplanes $\{H_{\mathcal{P}}\}_{\mathcal{P}\in\Pi}$ form a maximal tree $T$, and the edges in $\mathbb{S}_{\Gamma}^{\Pi}\setminus T$ are oriented and labelled by $V(\Gamma)$. Collapsing the maximal tree $T$ yields a rose with $n$ petals, each labelled by a generator of $A_{\Gamma}\cong F_{n}$ – this is a copy of the Salvetti complex.
@@
    The blowup $\mathbb{S}_{\Gamma}^{\mathcal{P}}$ of the Salvetti complex by a single partition $\mathcal{P}$ has the following structure. It has two vertices, say $x_{1}$ and $x_{2}$. The carrier $\kappa(H_{\mathcal{P}})$ is isomorphic to the product of an interval with the Salvetti complex for the full subgraph on $\operatorname{lk}({\mathcal{P}})$. Let $v\in V(\Gamma)$. If there are two edges dual to $H_{v}$, then $v\in\operatorname{lk}({\mathcal{P}})$. If there is only one edge dual to the (oriented) hyperplane $H_{v}$ and it terminates at $x_{i}$, then $v\in P_{i}$; if the edge originates at $x_{i}$ then $v^{-1}\in P_{i}$. In this way one can recover $\mathcal{P}=\left(P_{1}\;|\;P_{2}\;|\;\operatorname{lk}({\mathcal{P}})\right)$ from the blowup $\mathbb{S}_{\Gamma}^{\mathcal{P}}$.
@@
If $\alpha$ is an untwisted marking, we call the equivalence class of the pair $(X,\alpha)$ an untwisted marked $\Gamma$-complex. If $X\cong\mathbb{S}_{\Gamma}$, we call the equivalence class of $(X,\alpha)$ a marked Salvetti. In this paper, ‘marked $\Gamma$-complex’ will always mean ‘untwisted marked $\Gamma$-complex’.
@@
Suppose that $\Pi=\{\mathcal{P}\}$. For each $m\in\operatorname{max}({\mathcal{P}})$, there is a cube complex isomorphism $h_{m}$ of $\mathbb{S}_{\Gamma}^{\mathcal{P}}$ with the following property. Letting $c_{\mathcal{P}}^{-1}$ be a homotopy inverse for $c_{\mathcal{P}}$, the composition
$\smash{c_{\mathcal{P}}\circ h_{m}\circ c_{\mathcal{P}}^{-1}\colon\mathbb{S}_{\Gamma}\to\mathbb{S}_{\Gamma}^{\mathcal{P}}\xrightarrow{\cong}\mathbb{S}_{\Gamma}^{\mathcal{P}}\to\mathbb{S}_{\Gamma}}$
induces the $\Gamma$-Whitehead automorphism $\varphi(\mathcal{P},m)$ on $A_{\Gamma}=~\pi_{1}\left(\mathbb{S}_{\Gamma}\right)$.
An acyclic set $\mathcal{T}$ of hyperplanes in a $\Gamma$-complex is called treelike if the collapse $c_{\mathcal{T}}\colon X\to X\smash{\sslash}\mathcal{T}$ gives a cube complex isomorphic to $\mathbb{S}_{\Gamma}$.
@@
Hence the star of a marked Salvetti $(\mathbb{S}_{\Gamma},\alpha)$ is a cube complex, with one $k$-cube for each compatible set $\Pi$ such that $|\Pi|=k$. We will denote this cube by $c\left(\emptyset,\Pi;\alpha\right)$, which specifies its two extreme vertices and the marking with which we view it. Faces of this cube correspond to subsets $\Pi_{1}$$\Pi_{2}$ with $\emptyset\subseteq\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi$, and will be denoted $c\left(\Pi_{1},\Pi_{2};\alpha\right)$. We suppress the marking from the notation when it is not needed. Note that a cube $c\left(\Pi_{1},\Pi_{2}\right)$ may lie in the star of two different marked Salvettis $(\mathbb{S}_{\Gamma},\alpha)$ and $(\mathbb{S}_{\Gamma},\beta)$. In this case, writing $c\left(\Pi_{1},\Pi_{2};\alpha\right)$ simply conveys that we are presently thinking of this cube as a subface of a cube in the star of $(\mathbb{S}_{\Gamma},\alpha)$.
@@
Charney–Stambaugh–Vogtmann [CSV17] also build a full untwisted Outer space $\mathcal{O}_{\Gamma}^{U}$ for $U(A_{\Gamma})$, in analogy with Culler–Vogtmann Outer space $CV_{n}$ for the free group $F_{n}$. The spine $K_{\Gamma}$ naturally sits inside $\mathcal{O}_{\Gamma}^{U}$ as a deformation retract, generalising the way $K_{n}$, the spine of $CV_{n}$, sits inside reduced Culler–Vogtmann Outer space $CV_{n}^{red}$ (the deformation retract of $CV_{n}$ obtained by restricting only to graphs with no separating edges). However, we will only be interested in the intrinsic structure of $K_{\Gamma}$ as the geometric realisation of a poset.
@@
The geometric control afforded us by Theorem 2.7 yields the immediate corollary